the measure of one interior angle of a regular polygon is 6 greater than twice the measure of an exterior angle at the same vertex. How many sides does the polygon have?
Answer by stanbon(75887)

the measure of one interior angle of a regular polygon is 6 greater than twice the measure of an exterior angle at the same vertex. How many sides does the polygon have?
---------
Each int/ext pair of angles is supplementary.
Equation:
int + ext = 180
int = 2*ext + 6
------
Substitute for int and solve for ext:
2ext + 6 + ext = 180
3ext = 174
ext = 58 degrees
==========
The sum of all of the exterior angle is 360 degrees
# of sides = # of ext angles = 360/58 = 6.2069
----------
Note: Part of the directions for the problem must be incorrect.
The # of sides should be a whole number.
==============
Cheers,
Stan H.